Three loaves of bread and two jars of peanut butter cost cost $5.49. Four loaves of bread and three jars of peanut butter cost $7.75. Find the rice of a loaf of bread and a jar of peanut butter.
Found 2 solutions by solver91311, stanbon:
Answer by solver91311(24713)







Solve the system. is price of the bread, and is the price of the peanut butter.

John


Answer by stanbon(75887)

Three loaves of bread and two jars of peanut butter cost cost $5.49.
3b + 2p = 5.49
------------------------
Four loaves of bread and three jars of peanut butter cost $7.75
4b + 3p = 7.75
------------------------
Find the price of a loaf of bread and a jar of peanut butter.
------
Multiply thru 1st by 3
Multiply thru 2nd by 2
------------------------------
9b + 6p = 3*5.49
8b + 6p = 2*7.75
---------------------
Subtract 2nd from 1st and solve for "b":
b = $0.97 (price of loaf of bread)
-----------------
Substitute into 4b+3p = 7.75 to solve for "p":
4(0.97) + 3p = 7.75
3p = 3.87
p = $1.29 (price of peanut butter)
==================================
Cheers,
Stan H.
==================================